A road has a hill with a top in the shape of a circular arc of radius 32.0 m. How fast can a car go over the top of the hill without losing contact with the ground?
Answer #1

Solution) Radius r = 32 m

Speed V = ?

For circular path it is centripetal force Fc that acts

F = Fc

ma = (m(V^2))/(r)

a = (V^2)/(r)

V^2 = ar

V = (ar)^(1/2)

Here a = g = 9.8 m/s^2

V = (9.8×32)^(1/2)

V = 17.70 m/s